How many kilocalories are generated when the brakes are used to bring a 1100-kg car to rest from a speed of 85km/h ? 1 kcal = 4186 J
Answer #1

speed v =85 km/h =85*5/18 = 23.61 m/s

energy =0.5m(v^2-u^2) =0.5*1100(23.612 -0) =306616.5 J

energy E =306616.5/4186 = 73.24 kcal
Answer #2
Here you should use the kinetic energy formula